# Optimal Selection Using Algorithmic Rankings with Side Information

### v2

@@ D.2 Beyond-the-superstar strategy
$\prod_{i\in[N]\mid P[i\in\sigma_{1:k}]^{\prime}>0}(1-p_{i})\cdot P[i\in\sigma_{1:k}]^{\prime}\prod_{i\in[N]\mid P[i\in\sigma_{1:k}]>0}(1-p_{i})\cdot P[i\in\sigma_{1:k}]$
@@
which is higher, so the probability of the top candidate being picked only increases. Next, we consider the case where the window stays the same but the accuracy increases. By Definition 4 and our reasoning in the proof of Theorem 5 we know that increasing the accuracy of the ranking tool only increases the total probability of the top $k$ elements being busy. Thurs, this means that the first term increases and the second term increases - whether or not the total term increases or decreases depends on the relative change.
@@
See 7 Note that if we are in the superstar setting, there exists only a single pair of unique candidate types to compare, $v_{1}>v_{2}$, so all of the “votes” will go in the same direction. First, we will consider the setting where the first candidate is busy.
@@
Note that the LHS is equal to $1/R$ as calculated in Algorithm 1. For the $RHS$, $\sigma$ is defined as a permutation with the high value candidate ranked 11st, and $\tilde{}\sigma$ is a permutation with it ranked in index $j$. Note that $P[\sigma]\neq P[\sigma^{1}]$, because the latter is the total probability that the high value candidate is ranked 1st, while $P[\sigma]$ is the probability of a single permutation occurring where the high value candidate is ranked first. However, $P[\sigma^{1}]=N\cdot P[\sigma]$ and $P[\sigma^{j}]=N\cdot P[\tilde{}\sigma]$, and so $P[\sigma]/P[\tilde{}\sigma]=P[\sigma^{1}]/P[\sigma^{j}]$, as desired. Similar analysis holds for the setting where the first candidate is free.
@@
We will prove these properties in sequence. First, we note that 16 gives us a closed-form solution for the relative probability of finding a permutation with two items flipped:
@@
$\tilde{G}_{ik}=r_{i,k}\cdot P[\sigma^{\prime}]\cdot(v_{k}/\gamma-v_{i})+P[\tilde{}\sigma^{\prime}]\cdot(v_{i}/\gamma-v_{k})$
@@
  • If $C_{1},C_{2}>0$, then $\tilde{G}_{ik}$ can be written as $r_{ik}\cdot P[\sigma^{\prime}]\cdot(-C_{1})+P[\tilde{}\sigma^{\prime}]\cdot(-C_{2})$, and thus is negative.
@@
  • If $C_{1}<0,C_{2}>0$, then we know because $G_{ik}>0$ we have $r_{ik}\cdot C_{1}<C_{2}$, which implies that $C_{1}>-C_{2}$, and thus $C_{1}\cdot r_{ik}>C_{2}$, which implies that $\tilde{G}_{ik}<0$
@@
    which is satisfied by assumption
@@
$0\leq\frac{1}{P[s]}\sum_{\sigma\in\mathcal{P},v_{\sigma_{1}}>v_{\sigma_{j}}}P[s\mid\sigma]\cdot\left(r_{\sigma_{1},\sigma_{j}}\cdot P[\sigma]\cdot(v_{\sigma_{j}}/\gamma_{\sigma_{j}}-v_{\sigma_{1}})+P[\tilde{}\sigma]\cdot(v_{\sigma_{1}}/\gamma_{\sigma_{1}}-v_{\sigma_{j}})\right)$ (6)
@@
$0\leq\frac{1}{P[s]}\sum_{i,k\mid v_{i}>v_{k}}\left(\left(\sum_{\sigma\mid\sigma_{1}=i,\sigma_{j}=k}P[s\mid\sigma]\cdot P[\sigma]\right)\cdot\left(r_{\sigma_{1},\sigma_{j}}\cdot(v_{\sigma_{j}}/\gamma_{\sigma_{j}}-v_{\sigma_{1}})\right)+\left(\sum_{\tilde{}\sigma\mid\tilde{}\sigma_{1}=k,\tilde{}\sigma_{j}=i}P[s\mid\sigma]\cdot P[\tilde{}\sigma]\right)\cdot(v_{\sigma_{1}}/\gamma_{\sigma_{1}}-v_{\sigma_{j}})\right)$ (7)
@@
$G_{ik}=r_{i,k}\cdot P[\sigma^{\prime}]\cdot(v_{k}/\gamma_{k}-v_{i})+P[\tilde{}\sigma^{\prime}]\cdot(v_{i}/\gamma_{i}-v_{k})$
@@
$\frac{P[s\mid\sigma]}{P[s]}\sum_{\sigma,\tilde{}\sigma\in\mathcal{P}_{i,k,j}}-\left(r_{\sigma_{1},\sigma_{j}}\cdot P[\sigma]\cdot(v_{\sigma_{j}}/\gamma_{\sigma_{j}}-v_{\sigma_{1}})+P[\tilde{}\sigma]\cdot(v_{\sigma_{1}}/\gamma_{\sigma_{1}}-v_{\sigma_{j}})\right)\leq\max{\sigma,\rho}\frac{P[s\mid\sigma]}{P[s\mid\rho]}\cdot P[\mathcal{P}_{i,k,j}]\cdot(v_{i}-v_{k}/\gamma)$
@@
$P[\mathcal{P}_{i,k,j}]\cdot(v_{i}/\gamma-v_{k})$

### v3

@@ D.2 Beyond-the-superstar strategy
$\prod_{i\in[N]\mid P[i\in\sigma_{1:k}]^{\prime}>0}(1-p_{i})\cdot P[i\in\sigma_{1:k}]^{\prime}\geq\prod_{i\in[N]\mid P[i\in\sigma_{1:k}]>0}(1-p_{i})\cdot P[i\in\sigma_{1:k}]$
@@
which is higher, so the probability of the top candidate being picked only increases. Next, we consider the case where the window stays the same but the accuracy increases. By Definition 4 and our reasoning in the proof of Theorem 5 we know that increasing the accuracy of the ranking tool only increases the total probability of the top $k$ elements being busy. Thus, this means that the first term increases and the second term increases - whether or not the total term increases or decreases depends on the relative change.
@@
See 7
Proof.
Note that if we are in the superstar setting, there exists only a single pair of unique candidate types to compare, $v_{1}>v_{2}$, so all of the “votes” will go in the same direction. First, we will consider the setting where the first candidate is busy.
@@
Note that the LHS is equal to $1/R$ as calculated in Algorithm 1. For the $RHS$, $\sigma$ is defined as a permutation with the high value candidate ranked 11st, and $\tilde{}\sigma$ is a permutation with it ranked in index $j$. Note that $P[\sigma]\neq P[\sigma^{1}]$, because the latter is the total probability that the high value candidate is ranked 1st, while $P[\sigma]$ is the probability of a single permutation occurring where the high value candidate is ranked first. However, $P[\sigma^{1}]=N\cdot P[\sigma]$ and $P[\sigma^{j}]=N\cdot P[\tilde{}\sigma]$, and so $P[\sigma]/P[\tilde{}\sigma]=P[\sigma^{1}]/P[\sigma^{j}]$, as desired. Similar analysis holds for the setting where the first candidate is free. ∎
@@
We will prove these properties in sequence. First, we note that Lemma 14 gives us a closed-form solution for the relative probability of finding a permutation with two items flipped:
@@
$\tilde{G}_{ik}=P[\sigma^{\prime}]\cdot(v_{k}/\gamma-v_{i})+r_{i,k}\cdot P[\tilde{}\sigma^{\prime}]\cdot(v_{i}/\gamma-v_{k})$
@@
  • If $C_{1},C_{2}>0$, then $\tilde{G}_{ik}$ can be written as $P[\sigma^{\prime}]\cdot(-C_{1})+r_{ik}\cdot P[\tilde{}\sigma^{\prime}]\cdot(-C_{2})$, and thus is negative.
@@
  • If $C_{1}<0,C_{2}>0$, then we know because $G_{ik}>0$
    $0<G_{ik}\leq r_{ik}\cdot C_{1}+C_{2}$
    we have $\left|r_{ik}\cdot C_{1}\right|<C_{2}$, which implies that
    $\tilde{G}_{ik}\leq-C_{2}-r\cdot C_{1}\leq 0$
@@
    which is satisfied by assumption.
@@
$0\leq\frac{1}{P[s]}\sum_{\sigma\in\mathcal{P},v_{\sigma_{1}}>v_{\sigma_{j}}}P[s\mid\sigma]\cdot\left(P[\sigma]\cdot(v_{\sigma_{j}}/\gamma_{\sigma_{j}}-v_{\sigma_{1}})+r_{\sigma_{1},\sigma_{j}}\cdot P[\tilde{}\sigma]\cdot(v_{\sigma_{1}}/\gamma_{\sigma_{1}}-v_{\sigma_{j}})\right)$ (6)
@@
$0\leq\frac{1}{P[s]}\sum_{i,k\mid v_{i}>v_{k}}\left(\left(\sum_{\sigma\mid\sigma_{1}=i,\sigma_{j}=k}P[s\mid\sigma]\cdot P[\sigma]\right)\cdot\left((v_{\sigma_{j}}/\gamma_{\sigma_{j}}-v_{\sigma_{1}})\right)+\left(\sum_{\tilde{}\sigma\mid\tilde{}\sigma_{1}=k,\tilde{}\sigma_{j}=i}P[s\mid\sigma]\cdot P[\tilde{}\sigma]\right)\cdot r_{\sigma_{1},\sigma_{j}}\cdot(v_{\sigma_{1}}/\gamma_{\sigma_{1}}-v_{\sigma_{j}})\right)$ (7)
@@
$G_{ik}=P[\sigma^{\prime}]\cdot(v_{k}/\gamma_{k}-v_{i})+r_{i,k}\cdot P[\tilde{}\sigma^{\prime}]\cdot(v_{i}/\gamma_{i}-v_{k})$
@@
$\frac{P[s\mid\sigma]}{P[s]}\sum_{\sigma,\tilde{}\sigma\in\mathcal{P}_{i,k,j}}-\left(P[\sigma]\cdot(v_{\sigma_{j}}/\gamma_{\sigma_{j}}-v_{\sigma_{1}})+r_{\sigma_{1},\sigma_{j}}\cdot P[\tilde{}\sigma]\cdot(v_{\sigma_{1}}/\gamma_{\sigma_{1}}-v_{\sigma_{j}})\right)\leq\max{\sigma,\rho}\frac{P[s\mid\sigma]}{P[s\mid\rho]}\cdot P[\mathcal{P}_{i,k,j}]\cdot(v_{i}-v_{k}/\gamma)$
@@
$P[\mathcal{P}_{i,k,j}]\cdot r_{1,k}\cdot(v_{i}/\gamma-v_{k})$